# Random Forests as Statistical Procedures: Design, Variance, and Dependence

### v1

@@ Abstract
Random forests are widely used prediction procedures, yet are typically described algorithmically rather than as statistical designs acting on a fixed dataset. We develop a finite‑sample, design‑based formulation of random forests in which each tree is an explicit randomized conditional regression function. This perspective yields an exact variance identity for the forest predictor that separates finite‑aggregation variability from a structural dependence term that persists even under infinite aggregation. We further decompose both single‑tree dispersion and inter‑tree covariance using the laws of total variance and covariance, isolating two fundamental design mechanisms—reuse of training observations and alignment of data‑adaptive partitions. These mechanisms induce a strict covariance floor, demonstrating that predictive variability cannot be eliminated by increasing the number of trees alone. The resulting framework clarifies how resampling, feature‑level randomization, and split selection govern resolution, tree variability, and dependence, and establishes random forests as explicit finite‑sample statistical designs whose behavior is determined by their underlying randomized construction.
@@ 1.2 A design-based perspective
Recent work has broadly synthesized how Random Forests work as adaptive smoothers Curth et al. (2024). In this paper, we approach Random Forests through a difference lens – we define random forests as finite-sample statistical procedures generated by an explicit randomized design acting on a fixed observed dataset $\mathcal{D}_{n}=\{(X_{i},Y_{i})\}_{i=1}^{n}$. Each tree is constructed by randomized operations (e.g. observation selection (Breiman, 1996), random subspace selection for candidate splits (Ho, 1998), and stochastic split choice), that determine how the observed outcomes are combined to form predictions. All variability in the forest predictor is induced by this tree-generating design; the data $\mathcal{D}_{n}$ are treated as fixed and no probability model is imposed on the covariates or outcomes. All results are exact at fixed $n$.
When the dataset itself is viewed as random under repeated sampling, the unconditional variability of the forest predictor decomposes into a sampling component and a design component. The sampling component reflects instability of the induced infinite-forest target (the deterministic limit of the forest predictor as the number of trees goes to infinity) across datasets and, under standard stability conditions, decreases at the $O(n^{-1})$ rate. By contrast, the design component captures finite-sample variability introduced by the forest’s randomized construction acting on a fixed dataset and depends explicitly on aggregation, subsampling, and partitioning choices. Section 2 formalizes this decomposition and motivates our focus on the design-induced variability that distinguishes random forests as statistical procedures.
@@ 2.1 Variance decomposition and scope
We distinguish two sources of variability in random forest prediction: algorithmic variability induced by the forest-generating randomization acting on a fixed realized dataset, and sampling variability arising from repeated draws of the dataset from an underlying population.
Let $\mathcal{D}_{n}^{\star}$ denote a generic sample of size $n$ from the population, and let $\hat{f}_{B}(x;\mathcal{D},\theta)$ denote the $B$-tree forest predictor constructed from dataset $\mathcal{D}$ using design randomization $\theta$. The unconditional variance across repeated datasets and forest constructions decomposes by the law of total variance as
$\operatorname{Var}\!\bigl(\hat{f}_{B}(x;\mathcal{D}_{n}^{\star},\theta)\bigr)=\mathbb{E}\!\left[\operatorname{Var}\!\bigl(\hat{f}_{B}(x;\mathcal{D}_{n}^{\star},\theta)\mid\mathcal{D}_{n}^{\star}\bigr)\right]+\operatorname{Var}\!\left(\mathbb{E}\!\bigl[\hat{f}_{B}(x;\mathcal{D}_{n}^{\star},\theta)\mid\mathcal{D}_{n}^{\star}\bigr]\right).$
This decomposition is standard – see for example a similar representatin in (Hastie et al., 2009) and (Curth et al., 2024) – where it is used to separate sampling variance from algorithmic variance in the convetional bias-variance framework. Our use of this identity differs in purpose. We take the next step and condition on the realized dataset $D_{n}^{\star}$. In doing so, we isolate the first term as the design-induced variablity arising from the random forest applied to a fixed dataset, while the second captures sampling variability across repeated datasets. Under standard stability conditions, the sampling component is generically $O(n^{-1})$, reflecting conventional large-sample behavior; a derivation based on asymptotic linearity is given in Appendix A, with related results in Mentch and Hooker (2016); Wager and Athey (2018).
Unlike sampling variability, the design-induced variability persists at finite sample sizes and depends explicitly on forest hyperparameters governing aggregation, subsampling, and partitioning resolution. These choices control single-tree variability, the strength of dependence between trees, and the extent to which aggregation can reduce variance. Accordingly, for the remainder of this paper we fix the realized dataset $\mathcal{D}_{n}$ and focus on an exact finite-sample analysis of the design-induced variability $\operatorname{Var}(\hat{f}_{B}(x;\mathcal{D}_{n},\theta))$. All probabilistic quantities are therefore evaluated under the design-induced distribution conditional on a fixed $\mathcal{D}_{n}$; for notational simplicity we suppress the subscript $D_{n}$ on $\operatorname{Var}(\cdot)$, $\operatorname{Cov}(\cdot)$, and $E(\cdot)$.
@@ 2.2 Tree-level randomized regression functions
We formalize a single tree as a randomized prediction rule acting on a fixed observed dataset $\mathcal{D}_{n}=\{(X_{i},Y_{i})\}_{i=1}^{n}$, with $X_{i}\in\mathcal{X}\subset\mathbb{R}^{p}$ and $Y_{i}\in\mathbb{R}$. A tree is generated by applying a randomized recursive partitioning algorithm to $\mathcal{D}_{n}$, incorporating operations such as observation selection, random subspace selection for candidate splits, and stochastic split choice. A particular realization of this algorithmic randomization is indexed by $\theta$, and the resulting predictor is denoted by $T_{\theta}$. For fixed $\theta$, the tree and its induced prediction rule are deterministic functions of $\mathcal{D}_{n}$.
@@
The weights $W_{i}(x;\theta)$ are normalized membership indicators derived entirely from the averaging set $A_{\theta}(x)$ and introduce no additional randomness beyond the grouping induced by the tree. As $\theta$ varies under the tree-generating mechanism, the prediction $T_{\theta}(x)$ varies solely through the induced averaging set, equivalently through the collection of indicators $\{M_{i,\theta}(x)\}_{i=1}^{n}$. Although the observed outcomes $Y_{1:n}$ are treated as fixed, $A_{\theta}(x)$ depends on them because split selection optimizes impurity criteria computed from the resampled outcomes determined by $\theta$.
Representations of random forests as adaptive weighted averages of the training responses are well established (Lin and Jeon, 2006; Biau and Devroye, 2010; Scornet et al., 2015) and underlie extensions such as quantile regression forests (Meinshausen, 2006). Our contribution is not the algebraic form itself, but the probabilistic framing: we treat the random weight vector $W(x;\theta)$ as the primary random object in the design-based probability space. In this view, the infinite-aggregation predictor arises as the expectation of this random weight vector, and properties such as resolution and structural dependence are governed by its distribution rather than by the fitted tree viewed as a fixed estimator.
@@
Figure 1: Random forests as randomized local averaging on fixed outcomes. $\mathcal{D}_{n}$ are fixed. Independent draws of the tree-generating design $\theta_{1},\theta_{2},\theta_{3}$ induce realized tree structures and tree specific terminal-node membership sets $A_{\theta_{b}}(x)$ for prediction point $x$. Each tree prediction is the average over indexed outcomes in its membership set, and averaging over independent draws yields $\hat{f}_{B}(x)$ and $f_{\infty}(x)$.
@@ 2.3 The forest predictor and its design-based target
The aggregation level $B$ controls only the amount of Monte Carlo averaging and does not alter the distribution of the underlying tree-level rule. Let $\theta$ denote a generic draw from the tree-generating mechanism and define the infinite-aggregation target
$f_{\infty}(x)=\mathbb{E}\!\left[T_{\theta}(x)\right],$
where the expectation is taken with respect to the tree-generating randomization acting on the fixed dataset $\mathcal{D}_{n}$. This quantity is fully determined by the forest design and the observed data and represents the deterministic prediction induced by the procedure under infinite aggregation. The finite forest predictor $\hat{f}_{B}(x)$ is a Monte Carlo approximation to $f_{\infty}(x)$, satisfying $\mathbb{E}[\hat{f}_{B}(x)]=f_{\infty}(x)$. Finite forests fluctuate around this target due to limited aggregation, while structural dependence between tree-level prediction rules induces a non-zero covariance component as $B\to\infty$. The variance analysis that follows makes this separation explicit.
@@ 3.1 Finite-sample variance identity
We now characterize the finite-sample variance of the random forest predictor $\hat{f}_{B}(x)$ under the design-induced distribution. Recall that $\hat{f}_{B}(x)$ is the average of $B$ independent realizations $T_{\theta_{b}}(x)$ of the randomized tree-level prediction rule $T_{\theta}(x)$ acting on the fixed dataset $\mathcal{D}_{n}$. Define
$\sigma_{T}^{2}(x)=\operatorname{Var}\!\bigl(T_{\theta}(x)\bigr),\qquad C_{T}(x)=\operatorname{Cov}\!\bigl(T_{\theta}(x),T_{\theta^{\prime}}(x)\bigr),$
@@
$\operatorname{Var}\!\left(\hat{f}_{B}(x)\right)=\frac{1}{B}\,\sigma_{T}^{2}(x)+\frac{B-1}{B}\,C_{T}(x).$
@@
$\operatorname{Var}\!\bigl(\hat{f}_{B}(x)\bigr)=\frac{1}{B^{2}}\operatorname{Var}\!\left(\sum_{b=1}^{B}T_{\theta_{b}}(x)\right)=\frac{1}{B^{2}}\sum_{b=1}^{B}\sum_{b^{\prime}=1}^{B}\operatorname{Cov}\!\bigl(T_{\theta_{b}}(x),T_{\theta_{b^{\prime}}}(x)\bigr).$
By exchangeability of the tree-generating mechanism, $\operatorname{Cov}(T_{\theta_{b}}(x),T_{\theta_{b}}(x))=\sigma_{T}^{2}(x)$ for all $b$ and $\operatorname{Cov}(T_{\theta_{b}}(x),T_{\theta_{b^{\prime}}}(x))=C_{T}(x)$ for all $b\neq b^{\prime}$. Therefore
$\operatorname{Var}\!\bigl(\hat{f}_{B}(x)\bigr)=\frac{1}{B^{2}}\Bigl(B\,\sigma_{T}^{2}(x)+B(B-1)\,C_{T}(x)\Bigr)=\frac{1}{B}\sigma_{T}^{2}(x)+\frac{B-1}{B}C_{T}(x).$
@@ 3.2 Decomposing the single-tree variance
We now decompose the single-tree variance term $\sigma_{T}^{2}(x)$ appearing in Theorem 1. Let $\mathcal{I}_{\theta}=(I_{\theta,1},\dots,I_{\theta,n})$ denote the resampling indicators for the tree indexed by $\theta$, where $I_{\theta,i}=1$ if observation $i$ is exposed to the tree construction and $I_{\theta,i}=0$ otherwise. For a fixed prediction point $x$, the resampling indicators determine which observations are eligible to appear in the random averaging set $A_{\theta}(x)$, while the remaining algorithmic randomness—such as feature subsampling and split selection—determines how $x$ is grouped with those observations. Conditioning on $\mathcal{I}_{\theta}$ and applying the law of total variance yields
$\sigma_{T}^{2}(x)=\mathbb{E}\!\left[\operatorname{Var}\!\left(T_{\theta}(x)\mid\mathcal{I}_{\theta}\right)\right]+\operatorname{Var}\!\left(\mathbb{E}\!\left[T_{\theta}(x)\mid\mathcal{I}_{\theta}\right]\right).$ (1)
The first term in (1) captures variability arising from random splitting decisions after the resampling pattern is fixed, while the second captures variability induced by changes in the resampling realization itself, which alter which observations are eligible to be grouped with $x$. We now focus on the first component and examine how randomness in local grouping contributes to single-tree variability.
@@ 3.3 Decomposing the within-resample variance term
$V_{\mathrm{in}}(x)=\mathbb{E}\!\left[\operatorname{Var}\!\left(T_{\theta}(x)\mid\mathcal{I}_{\theta}\right)\right].$
@@
Conceptually, $V_{\mathrm{in}}(x)$ measures how sensitive the prediction at $x$ is to the particular way the tree is grown, holding fixed which observations are available to it. Even with the same resampled data, different choices of candidate split variables, cutpoints, or tie-breaking can route $x$ into different terminal nodes and therefore group it with different subsets of observations. Under a fixed resampling pattern, the remaining tree-level randomization affects $T_{\theta}(x)$ only through the induced averaging set $A_{\theta}(x)$. Conditioning on $A_{\theta}(x)$ therefore fixes the local grouping used to form the prediction at $x$. Applying the law of total variance with $A_{\theta}(x)$ as the intermediate conditioning object yields
@@
$V_{\mathrm{in}}(x)=\mathbb{E}\!\left[\operatorname{Var}\!\left(T_{\theta}(x)\mid\mathcal{I}_{\theta},A_{\theta}(x)\right)\right]+\mathbb{E}\!\left[\operatorname{Var}\!\left(\mathbb{E}\!\left[T_{\theta}(x)\mid\mathcal{I}_{\theta},A_{\theta}(x)\right]\mid\mathcal{I}_{\theta}\right)\right].$ (2)
@@
The first term in (2) does not represent additional algorithmic randomness. Once both the resampling pattern and the averaging set at $x$ are fixed, the tree prediction is deterministic. This component reflects how intrinsic outcome dispersion at the observations grouped with $x$ appears in the value of a fixed local average. Conditional on $(\mathcal{I}_{\theta},A_{\theta}(x))$,
$\operatorname{Var}\!\left(T_{\theta}(x)\mid\mathcal{I}_{\theta},A_{\theta}(x)\right)=\sum_{i=1}^{n}W_{i}(x;\theta)^{2}\,\operatorname{Var}(Y_{i}\mid\mathcal{I}_{\theta},A_{\theta}(x))\;+\;\mathcal{C}_{\theta}(x),$
where $\mathcal{C}_{\theta}(x)$ collects cross-unit covariance terms and is zero under conditional independence across observational units. Conditioning on $(\mathcal{I}_{\theta},A_{\theta}(x))$ fixes only how observations are grouped with $x$ and does not alter their outcome-generating behavior, so $\operatorname{Var}(Y_{i}\mid\mathcal{I}_{\theta},A_{\theta}(x))=\operatorname{Var}(Y\mid X=X_{i})$. Writing this deterministic quantity as $\sigma_{i}^{2}$,
$\mathbb{E}\!\left[\operatorname{Var}\!\left(T_{\theta}(x)\mid\mathcal{I}_{\theta},A_{\theta}(x)\right)\right]=\sum_{i=1}^{n}\sigma_{i}^{2}\,\mathbb{E}\!\left[W_{i}(x;\theta)^{2}\right].$
This term captures the intrinsic statistical variability of the averaging rule itself: how outcome noise at contributing observations propagates through a fixed local average. The quantity $\sigma_{i}^{2}=\operatorname{Var}(Y\mid X=X_{i})$ is a population-level conditional second moment that characterizes the intrinsic variability of outcomes at the covariate location $X_{i}$. It is not a parameter of the randomized design and is not a source of variability under the design-based distribution; rather, it enters as a fixed scalar when the design-induced variance is expanded at a known averaging set. In this sense, $\sigma_{i}^{2}$ plays the same role as the observed outcomes $Y_{i}$ themselves: it is a fixed attribute of the statistical environment in which the design operates. The decomposition holds for any values of $\sigma_{i}^{2}$ and does not require their estimation.
@@
The second term in (2) captures variability arising purely from randomness in the induced averaging set. Even with the resampling pattern fixed, different realizations of the remaining tree-level randomization can induce different sets $A_{\theta}(x)$ and hence different local groupings of observations around $x$. Because $\mathbb{E}[T_{\theta}(x)\mid\mathcal{I}_{\theta},A_{\theta}(x)]=\sum_{i=1}^{n}W_{i}(x;\theta)\,Y_{i}$, where the weights are deterministic functions of the averaging set, this component measures variability in the prediction at $x$ induced by randomness in the grouping itself, holding the observed outcomes fixed:
$\mathbb{E}\!\left[\operatorname{Var}\!\left(\sum_{i=1}^{n}W_{i}(x;\theta)\,Y_{i}\;\middle|\;\mathcal{I}_{\theta}\right)\right].$
@@
It is zero if and only if the averaging set $A_{\theta}(x)$ is almost surely constant given $\mathcal{I}_{\theta}$. This term therefore reflects instability in how the tree-growing procedure defines local neighborhoods around $x$: different partitions lead to different groupings and hence to different predictions, even when the available data are the same.

### v2

@@ Abstract
Random forests are widely used prediction procedures, yet are typically described algorithmically rather than as statistical designs acting on a fixed set of covariates. We develop a finite‑sample, design‑based formulation of random forests in which each tree is an explicit randomized conditional regression function. This perspective yields an exact variance identity for the forest predictor that separates finite‑aggregation variability from a structural dependence term that persists even under infinite aggregation. We further decompose both single‑tree dispersion and inter‑tree covariance using the laws of total variance and covariance, isolating two fundamental design mechanisms—reuse of training observations and alignment of data‑adaptive partitions. These mechanisms induce a strict covariance floor, demonstrating that predictive variability cannot be eliminated by increasing the number of trees alone. The resulting framework clarifies how resampling, feature‑level randomization, and split selection govern resolution, tree variability, and dependence, and establishes random forests as explicit finite‑sample statistical designs whose behavior is determined by their underlying randomized construction.
@@ 1.2 A design-based perspective
Recent work has synthesized Random Forests as adaptive smoothing procedures (Curth et al., 2024). We adopt a complementary perspective. We treat a random forest as a finite-sample statistical procedure generated by an explicit randomized design acting on a realized covariate configuration. Specifically, we fix the observed covariate points $X$ and analyze predictive variability under the joint randomness of the outcome realization $(Y\mid X)$ and the tree-generating mechanism $\theta$.
This conditioning choice is deliberate. Fixing $X$ but not $Y$ preserves both sources of variability that affect the forest predictor at the moment it is constructed: the intrinsic dispersion of outcomes at the realized design points, and the algorithmic randomness introduced by the forest construction. A practitioner observes covariate profiles that are determined by the study design or sampling frame; what remains uncertain is both what outcomes those covariate profiles produce and how the tree-building algorithm partitions them. Conditioning on $X$ captures this joint uncertainty.
Within this framework, each tree is viewed as arising from randomized operations—such as observation selection (Breiman, 1996), random subspace selection for candidate splits (Ho, 1998), and stochastic split choice—that determine how the outcomes are locally averaged to form predictions. The tree-generating mechanism induces a probability distribution over terminal regions and weight vectors conditional on $X$, and all algorithmic variability is defined with respect to this design. Outcome randomness enters through the conventional fixed-design model $Y\mid X$. All results are finite-sample and hold at fixed $n$.
When the dataset itself is viewed as random under repeated sampling, the unconditional variability of the forest predictor decomposes into a sampling component and a design component. The sampling component reflects instability of the induced infinite-forest target (the deterministic limit of the forest predictor as the number of trees goes to infinity) across datasets and, under standard stability conditions, decreases at the $O(n^{-1})$ rate. By contrast, the design component captures variability introduced by the forest’s randomized construction acting on the realized covariate configuration and depends explicitly on aggregation, subsampling, and partitioning choices. Section 2 formalizes this decomposition and motivates our focus on the design-induced variability that distinguishes random forests as statistical procedures.
@@ 2.1 Variance decomposition and scope
We distinguish two sources of variability in random forest prediction: procedural variability arising from the forest-generating randomization acting on a realized covariate configuration, and sampling variability arising from repeated draws of the dataset from an underlying population.
Let $\mathcal{D}_{n}^{\star}=\{(X_{i},Y_{i})\}_{i=1}^{n}$ denote a generic sample of size $n$ from the population, and let $\hat{f}_{B}(x;\mathcal{D},\theta)$ denote the $B$-tree forest predictor constructed from dataset $\mathcal{D}$ using design randomization $\theta$. To isolate variability conditional on the realized covariate configuration, we apply the law of total variance with respect to $X$. This yields
$\operatorname{Var}\!\bigl(\hat{f}_{B}(x;\mathcal{D}_{n}^{\star},\theta)\bigr)=\mathbb{E}\!\left[\operatorname{Var}\!\bigl(\hat{f}_{B}(x;\mathcal{D}_{n}^{\star},\theta)\mid X\bigr)\right]+\operatorname{Var}\!\left(\mathbb{E}\!\bigl[\hat{f}_{B}(x;\mathcal{D}_{n}^{\star},\theta)\mid X\bigr]\right).$
The first term represents the finite-sample variability of a forest at a realized covariate configuration, where randomness arises jointly from outcome variation under the fixed-design model $Y\mid X$ and from the tree-generating mechanism $\theta$. The second term reflects sampling variability of the covariate configuration itself; it captures how the induced infinite-forest target changes across repeated draws of $X$ from the population. Under standard stability conditions, this sampling component is generically $O(n^{-1})$; see Appendix A and related results in Mentch and Hooker (2016); Wager and Athey (2018).
Accordingly, for the remainder of this paper we fix the realized covariate configuration $X$ and analyze the finite-sample variability of $\hat{f}_{B}(x)$ under the joint randomness of $(Y\mid X)$ and the forest-generating mechanism $\theta$. All probabilistic quantities are therefore evaluated conditional on $X$.
@@ 2.2 Tree-level randomized regression functions
We formalize a single tree as a randomized prediction rule acting on a realized covariate configuration $X=\{X_{i}\}_{i=1}^{n}$, with $X_{i}\in\mathcal{X}\subset\mathbb{R}^{p}$. Outcomes satisfy a fixed-design model $Y\mid X$, and randomness in the tree arises jointly from the outcome realization under this model and from the algorithmic construction. A tree is generated by applying a randomized recursive partitioning algorithm to the training sample $(X,Y)$, incorporating operations such as observation selection, random subspace selection for candidate splits, and stochastic split choice. A particular realization of this algorithmic randomization is indexed by $\theta$, and the resulting predictor is denoted by $T_{\theta}$. At a fixed realization $(Y,\theta)$ and fixed covariate configuration $X$, the tree and its induced prediction rule are deterministic functions.
@@
The weights $W_{i}(x;\theta)$ are normalized membership indicators derived entirely from the averaging set $A_{\theta}(x)$ and introduce no additional randomness beyond that induced by the tree-generating mechanism. As $\theta$ varies under the tree-generating mechanism, the prediction $T_{\theta}(x)$ varies solely through the induced averaging set, equivalently through the collection of indicators $\{M_{i,\theta}(x)\}_{i=1}^{n}$. Although the covariate configuration $X$ is fixed, the averaging set $A_{\theta}(x)$ depends on the realized outcomes $Y$ because split selection optimizes impurity criteria computed from the training responses available to tree $\theta$.
Representations of random forests as adaptive weighted averages of the training responses are well established (Lin and Jeon, 2006; Biau and Devroye, 2010; Scornet et al., 2015) and underlie extensions such as quantile regression forests (Meinshausen, 2006). Our contribution is not the algebraic form itself, but the probabilistic framing: we treat the random weight vector $W(x;\theta)$ as the primary random object in the induced probability space. In this view, the infinite-aggregation predictor arises as the expectation of this random weight vector under the tree-generating mechanism conditional on $X$, and subsequent variance properties are governed by its distribution.
@@
Figure 1: Random forests as randomized local averaging on fixed outcomes. Independent draws of the tree-generating design $\theta_{1},\theta_{2},\theta_{3}$ induce realized tree structures and tree specific terminal-node membership sets $A_{\theta_{b}}(x)$ for prediction point $x$. Each tree prediction is the average over indexed outcomes in its membership set, and averaging over independent draws yields $\hat{f}_{B}(x)$ and $f_{\infty}(x)$.
@@ 2.3 The forest predictor and its design-based target
The aggregation level $B$ controls only the amount of Monte Carlo averaging and does not alter the distribution of the underlying tree-level rule. Let $\theta$ denote a generic draw from the tree-generating mechanism and define the infinite-aggregation target as the $\theta$-average conditional on the realized covariate configuration,
$f_{\infty}(x):=\mathbb{E}\!\left[T_{\theta}(x)\mid X\right].$
Because $X$ is fixed but the training outcomes remain random under the fixed-design model, $f_{\infty}(x)$ is random through $Y$. The finite forest predictor $\hat{f}_{B}(x)$ is a Monte Carlo approximation to $f_{\infty}(x)$ and satisfies $\mathbb{E}[\hat{f}_{B}(x)\mid X]=f_{\infty}(x)$. Finite forests fluctuate around this target due to limited aggregation, while structural dependence between tree-level prediction rules induces a non-vanishing covariance component as $B\to\infty$. The variance analysis that follows makes this separation explicit.
@@ 3.1 Finite-sample variance identity
We now characterize the finite-sample variance of the random forest predictor $\hat{f}_{B}(x)$ conditional on the realized covariate configuration $X$. Recall that $\hat{f}_{B}(x)$ is the average of $B$ independent realizations $T_{\theta_{b}}(x)$ of the randomized tree-level prediction rule $T_{\theta}(x)$, where randomness arises jointly from $(Y\mid X)$ and the tree-generating mechanism. Define
$\sigma_{T}^{2}(x)=\operatorname{Var}\!\bigl(T_{\theta}(x)\mid X\bigr),\qquad C_{T}(x)=\operatorname{Cov}\!\bigl(T_{\theta}(x),T_{\theta^{\prime}}(x)\mid X\bigr),$
@@
$\operatorname{Var}\!\left(\hat{f}_{B}(x)\mid X\right)=\frac{1}{B}\,\sigma_{T}^{2}(x)+\frac{B-1}{B}\,C_{T}(x).$
@@
$\operatorname{Var}\!\bigl(\hat{f}_{B}(x)\mid X\bigr)=\frac{1}{B^{2}}\operatorname{Var}\!\left(\sum_{b=1}^{B}T_{\theta_{b}}(x)\mid X\right)=\frac{1}{B^{2}}\sum_{b=1}^{B}\sum_{b^{\prime}=1}^{B}\operatorname{Cov}\!\bigl(T_{\theta_{b}}(x),T_{\theta_{b^{\prime}}}(x)\mid X\bigr).$
By exchangeability of the tree-generating mechanism, $\operatorname{Var}(T_{\theta_{b}}(x)\mid X)=\sigma_{T}^{2}(x)$ for all $b$ and $\operatorname{Cov}(T_{\theta_{b}}(x),T_{\theta_{b^{\prime}}}(x)\mid X)=C_{T}(x)$ for all $b\neq b^{\prime}$. Therefore
$\operatorname{Var}\!\bigl(\hat{f}_{B}(x)\mid X\bigr)=\frac{1}{B^{2}}\Bigl(B\,\sigma_{T}^{2}(x)+B(B-1)\,C_{T}(x)\Bigr)=\frac{1}{B}\sigma_{T}^{2}(x)+\frac{B-1}{B}C_{T}(x).$
@@ 3.2 Decomposing the single-tree variance
We now decompose the single-tree variance term $\sigma_{T}^{2}(x)=\operatorname{Var}(T_{\theta}(x)\mid X)$ appearing in Theorem 1. Let $\mathcal{I}_{\theta}=(I_{\theta,1},\dots,I_{\theta,n})$ denote the resampling indicators for the tree indexed by $\theta$, where $I_{\theta,i}=1$ if observation $i$ is exposed to the tree construction and $I_{\theta,i}=0$ otherwise. For a fixed prediction point $x$, the resampling indicators determine which observations are eligible to appear in the random averaging set $A_{\theta}(x)$, while the remaining algorithmic randomness—such as feature subsampling and split selection—determines how $x$ is grouped with those observations. Conditioning on $\mathcal{I}_{\theta}$ in addition to $X$ and applying the law of total variance yields
$\operatorname{Var}\!\bigl(T_{\theta}(x)\mid X\bigr)=\mathbb{E}\!\left[\operatorname{Var}\!\bigl(T_{\theta}(x)\mid\mathcal{I}_{\theta},X\bigr)\mid X\right]+\operatorname{Var}\!\left(\mathbb{E}\!\bigl[T_{\theta}(x)\mid\mathcal{I}_{\theta},X\bigr]\mid X\right).$ (1)
The first term in (1) captures variability arising from random splitting decisions after the resampling pattern is fixed, while the second captures variability induced by changes in the resampling realization itself, which alter which observations are eligible to be grouped with $x$. Both components are evaluated under the joint randomness of $(Y\mid X)$ and the tree-generating mechanism $\theta$, with $X$ treated as fixed throughout. We now focus on the first component and examine how randomness in local grouping contributes to single-tree variability.
@@ 3.3 Decomposing the within-resample variance term
$V_{\mathrm{in}}(x)=\mathbb{E}\!\left[\operatorname{Var}\!\left(T_{\theta}(x)\mid\mathcal{I}_{\theta},X\right)\;\middle|\;X\right].$
@@
Conceptually, $V_{\mathrm{in}}(x)$ measures how sensitive the prediction at $x$ is to the particular way the tree is grown, holding fixed which observations are available to it. Even with the same resampled data, different choices of candidate split variables, cutpoints, or tie-breaking can route $x$ into different terminal nodes and therefore group it with different subsets of observations. Under a fixed resampling pattern, the remaining tree-level randomization affects $T_{\theta}(x)$ only through the induced averaging set $A_{\theta}(x)$. Conditioning on $A_{\theta}(x)$ therefore fixes the local grouping used to form the prediction at $x$. Applying the law of total variance conditional on $X$, with $A_{\theta}(x)$ as the intermediate conditioning object, yields
@@
$V_{\mathrm{in}}(x)=\mathbb{E}\!\left[\operatorname{Var}\!\left(T_{\theta}(x)\mid\mathcal{I}_{\theta},A_{\theta}(x),X\right)\;\middle|\;X\right]+\mathbb{E}\!\left[\operatorname{Var}\!\left(\mathbb{E}\!\left[T_{\theta}(x)\mid\mathcal{I}_{\theta},A_{\theta}(x),X\right]\;\middle|\;\mathcal{I}_{\theta},X\right)\;\middle|\;X\right].$ (2)
@@
The first term in (2) does not represent additional algorithmic randomness. Once both the resampling pattern and the averaging set at $x$ are fixed, the remaining tree-level randomization has been removed; variability in this component arises only from outcome randomness under the fixed-design model, propagated through the local averaging rule determined by $(\mathcal{I}_{\theta},A_{\theta}(x),X)$. Conditional on $(\mathcal{I}_{\theta},A_{\theta}(x),X)$, the weights $W_{i}(x;\theta)$ are deterministic, and the tree prediction is a fixed linear combination of outcomes. Therefore
$\operatorname{Var}\!\left(T_{\theta}(x)\mid\mathcal{I}_{\theta},A_{\theta}(x),X\right)=\sum_{i=1}^{n}\sum_{j=1}^{n}W_{i}(x;\theta)\,W_{j}(x;\theta)\,\operatorname{Cov}(Y_{i},Y_{j}\mid A_{\theta}(x),X),$
which under conditional independence of outcomes across observational units reduces to
$\operatorname{Var}\!\left(T_{\theta}(x)\mid\mathcal{I}_{\theta},A_{\theta}(x),X\right)=\sum_{i=1}^{n}W_{i}(x;\theta)^{2}\,\operatorname{Var}(Y_{i}\mid A_{\theta}(x),X).$
Because $\mathcal{I}_{\theta}(x)$ is a component of the algorithmic randomization drawn independently of the outcome-generating process, it carries no information about $Y_{i}$ beyond $X$ and can be removed from the conditioning. The averaging set $A_{\theta}(x)$, by contrast, is determined jointly by the algorithmic randomization and the realized outcomes through data-adaptive split selection, and therefore remains in the conditioning. This is the variance of a weighted average of conditionally independent outcomes at a realized set of design points, with weights fixed by the tree structure. Taking expectations over the design yields
$\mathbb{E}\!\left[\operatorname{Var}\!\left(T_{\theta}(x)\mid\mathcal{I}_{\theta},A_{\theta}(x),X\right)\;\middle|\;X\right]=\mathbb{E}\!\left[\sum_{i=1}^{n}W_{i}(x;\theta)^{2}\,\operatorname{Var}(Y_{i}\mid A_{\theta}(x),X)\;\middle|\;X\right].$
This term captures the intrinsic statistical variability of the averaging rule: how outcome noise at the contributing observations propagates through a fixed local average. The conditional variances $\operatorname{Var}(Y_{i}\mid A_{\theta}(x),X)$ are properties of the outcome model at the realized design points, not parameters of the forest design. They enter as scaling factors that govern how much outcome dispersion is available for the weights to transmit. The design controls the weights; the outcome model controls the noise; this term measures their interaction at a realized grouping.
@@
The second term in (2) captures variability arising purely from randomness in the induced averaging set. Even with the resampling pattern fixed, different realizations of the remaining tree-level randomization can induce different sets $A_{\theta}(x)$ and hence different local groupings of observations around $x$. Because
$\mathbb{E}[T_{\theta}(x)\mid\mathcal{I}_{\theta},A_{\theta}(x),X]=\sum_{i=1}^{n}W_{i}(x;\theta)\,Y_{i},$
where the weights are deterministic functions of the averaging set, this component measures variability in the prediction at $x$ induced by randomness in the grouping itself:
$\mathbb{E}\!\left[\operatorname{Var}\!\left(\sum_{i=1}^{n}W_{i}(x;\theta)\,Y_{i}\;\middle|\;\mathcal{I}_{\theta},X\right)\;\middle|\;X\right].$
@@
It is zero if and only if the averaging set $A_{\theta}(x)$ is almost surely constant given $(\mathcal{I}_{\theta},X)$. This term therefore reflects instability in how the tree-growing procedure defines local neighborhoods around $x$: different partitions lead to different groupings and hence to different predictions, even when the available data are the same.